j=p/9+2 , p=6j+2 Values of p and j please?
Answer by TimothyLamb(4379)

---
j=p/9+2
p=6j+2
---
put the system of linear equations into standard form
---
x = j
y = p
---
x = y/9 + 2
y = 6x + 2
---
x - (1/9)y = 2
6x - y = -2
---
x - 0.111111y = 2
6x - y = -2
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = j = 6.6666667 = 6+2/3 = 20/3
y = p = 42
---
system matrix
1 -0.11111111
6 -1
---
inverse of system matrix
3 -0.33333333
18 -3
---
determinant of system matrix = -0.333333333333
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php